OpenStudy (anonymous):

Determine the equation of the axis of symmetry. 4x^2+6x-1

OpenStudy (campbell_st):

the axis of symmetry is found for a quadratic \[ax^2 + bx + c \] as \[x = \frac{-b}{2a}\] in your question a = 4 and b = 6

OpenStudy (anonymous):

would it be negative or positive? positive right?

OpenStudy (campbell_st):

you axis is \[x = \frac{-6}{2 \times 4}\] simplify it... and its negative.